A Complete Bibliography of Publications in ACM Letters on Programming Languages and Systems (LOPLAS)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.08

Title word cross-reference

77D [33].

90D [33].

Abstract [18]. after [15]. algorithm [29]. algorithmic [22]. alias [2]. analysis [2, 15, 36, 17, 18, 23, 39]. approach [18]. approximation [10]. array [9, 36]. assembly [13]. assignment [12]. assisted [24]. assumption [8].

bound [36].

C [35, 13]. call [17, 20]. case [3]. checks [9, 36]. closure [14]. code [16, 12, 6]. code-generator [16]. collection [24]. collector [5]. Coloring [1]. compilation [33]. Compiling [38, 13]. complexity [4]. computers [38]. computing [26]. conditions [7]. constant [41]. constraint [40]. context [32]. context-free [32]. control [26, 14]. conversion [37]. Cost [24]. Cost-effective [24]. CPS [15]. CPS-transformation [15].

dangling [35]. data [2, 11, 18, 42]. data-flow [2]. debugging [22]. Definitions [19]. dependence [26, 19, 14]. description [18]. design [11]. detection [35]. direct [37]. directed [21]. distance [19]. Distributed [20].

effective [24]. effects [3]. Efficient [17, 30, 2, 16, 39]. empirical [41]. Engineering [16]. errors [27]. estimating [4]. evaluation [21]. exclusion [29]. executable [28]. Explaining [27]. Extending [32].

floating [11]. floating-point [11]. flow [26, 2, 36]. form [12]. formalizations [7]. Fortran [33]. free [32].

garbage [5, 24]. Generalized [22]. Generating [6]. generator [31, 16]. global [8]. goal [21]. goal-directed [21]. grammars [32]. graph [17]. groundness [39].

hardware [24]. hardware-assisted [24].

imperative [18]. Implementing [29]. improving [18]. indirect [37]. information [2]. inline [3]. integration [2]. intermediate [21]. Interprocedural [41, 28, 2]. issues [7]. iterated [14].

join [14].

Lamport [29]. languages [27]. last [20]. latency [5]. linear [26]. locks [29]. logic [10, 39, 40, 20].

machines [6]. maintenance [21]. management [24]. memory [38]. message [38]. message-passing [38]. ML [13]. mobile [34]. multiple [29]. mutual [29].

near [26]. near-linear [26]. No [13]. nondeterministic [42].

object [24, 25]. object-oriented [25]. objects [34]. opacity [42]. Optimization [9, 18, 40, 20]. optimizer [12]. Optimizing [36]. oriented [25].

pairs [1]. parallel [6, 20]. passing [38]. path [10]. permutation [32]. phrases [32]. point [11]. pointer [18]. Polymorphic [4, 27]. portable [20]. Precise [28, 2, 39]. procedures [25]. program [4]. programming [20, 25]. programs [10, 35, 18, 39, 40]. propagation [41].

race [7]. range [9]. RE2C [31]. real [5, 24]. real-time [5, 24]. reasoning [8]. recursion [37]. reducible [26]. Reducing [5]. References [34, 35]. Referential [42]. refinement [42]. region [26]. regions [26]. register [1]. remote [34]. representation [30]. required [13].

scanner [31]. Semantics [40]. sets [30]. shared [38]. shared-memory [38]. Shortest [10]. side [3]. SIMD [6]. simple [16]. single [12]. slices [28]. Some [7]. space [24]. sparse [30]. standard [13]. static [23, 12]. strictness [15]. structures [18]. study [3, 41]. subscript [9]. substitution [3]. systems [4].

testing [22]. Thor [34]. time

[26, 4, 5, 24, 8]. transformation [15]. transitive [14]. type [27]. types [11].

Undecidability [23]. Unexpected [3]. Unified [33]. Unreachable [25]. Using [12, 29, 36].

values [21]. versatile [31].

References

Briggs:1992:CRP

[1] Preston Briggs, Keith D. Cooper, and Linda Torczon. Coloring register pairs. ACM Letters on Programming Languages and Systems, 1(1):3–13, March 1992. CO-DEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http:

//www.acm.org/pubs/toc/Abstracts/
1057-4514/130617.html.

Burke:1992:PEI

[2] Michael Burke and Jong-Deok Choi. Precise and efficient integration of interprocedural alias information into dataflow analysis. ACM Letters on Programming Languages and Systems, 1(1):14–21, March 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

URL http://www.acm.org/pubs/toc/
Abstracts/1057-4514/130618.html.

Cooper:1992:USE

[3] Keith D. Cooper, Mary W. Hall, and Linda Torczon. Unexpected side effects of inline substitution: a case study. ACM Letters on Programming Languages and Systems, 1(1):22–32, March 1992. CO-DEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http:

//www.acm.org/pubs/toc/Abstracts/
1057-4514/130619.html.

Dornic:1992:PTS

[4] Vincent Dornic, Pierre Jouvelot, and David K. Gifford. Polymorphic time systems for estimating program complexity. ACM Letters on Programming Languages and Systems, 1(1):33–45, March 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

URL http://www.acm.org/pubs/toc/
Abstracts/1057-4514/130620.html.

Johnson:1992:RLR

[5] Ralph E. Johnson. Reducing the latency of a real-time garbage collector. ACM Letters on Programming Languages and Systems, 1(1):46–58, March 1992. CO-DEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http:

//www.acm.org/pubs/toc/Abstracts/
1057-4514/130621.html.

McKenney:1992:GPC

[6] Bruce McKenney and Boleslaw K. Szymanski. Generating parallel code for SIMD machines. ACM Letters on Programming Languages and Systems, 1(1):59–73, March 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/1057-4514/
130622.html.

Netzer:1992:WRC

[7] Robert H. B. Netzer and Barton P. Miller. What are race conditions?: Some issues and formalizations. ACM Letters on Programming Languages and Systems, 1(1):74–88, March 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/1057-4514/
130623.html.

Singh:1992:RGT

[8] Ambuj K. Singh. On reasoning with the global time assumption. ACM Letters on Programming Languages and Systems, 1(1):89–103, March 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/1057-4514/
130624.html.

Asuru:1992:OAS

[9] Jonathan M. Asuru. Optimization of array subscript range checks. ACM Letters on Programming Languages and Systems, 1(2):109–118, June 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/1057-4514/
151392.html.

Dietrich:1992:SPA

[10] Suzanne W. Dietrich. Shortest path by approximation in logic programs. ACM Letters on Programming Languages and Systems, 1(2):119–137, June 1992. CO-DEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http:

//www.acm.org/pubs/toc/Abstracts/
1057-4514/151377.html.

Goldberg:1992:DFD

[11] David Goldberg. The design of floatingpoint data types. ACM Letters on Programming Languages and Systems, 1(2):138–151, June 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/1057-4514/
151373.html.

McConnell:1992:USS

[12] Carl McConnell and Ralph E. Johnson. Using static single assignment form in a code optimizer. ACM Letters on Programming Languages and Systems, 1(2):152–160, June 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/1057-4514/
151368.html.

Tarditi:1992:NAR

[13] David Tarditi, Peter Lee, and Anurag Acharya. No assembly required: compiling standard ML to C. ACM Letters on Programming Languages and Systems, 1(2):161–177, June 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/1057-4514/
151343.html.

Weiss:1992:TCC

[14] Michael Weiss. The transitive closure of control dependence: the iterated join. ACM Letters on Programming Languages and Systems, 1(2):178– 190, June 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

URL http://www.acm.org/pubs/toc/
Abstracts/1057-4514/151337.html.

Danvy:1992:CAS

[15] Olivier Danvy and John Hatcliff. CPStransformation after strictness analysis. ACM Letters on Programming Languages and Systems, 1(3):195–212, September 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

URL http://www.acm.org/pubs/toc/
Abstracts/1057-4514/151641.html.

Fraser:1992:ESE

[16] Christopher W. Fraser, David R. Hanson, and Todd A. Proebsting. Engineering a simple, efficient code-generator generator. ACM Letters on Programming Languages and Systems, 1(3):213– 226, September 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (elec-

tronic). URL http://storage.webhop.
net/documents/iburg.pdf;   http:/
/www.acm.org/pubs/toc/Abstracts/
1057-4514/151642.html;   http://
www.cs.princeton.edu/software/iburg/

.

Hall:1992:ECG

[17] Mary W. Hall and Ken Kennedy. Efficient call graph analysis. ACM Letters on Programming Languages and Systems, 1(3):227–242, September 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/1057-4514/
151643.html.

Hummel:1992:ADP

[18] Joseph Hummel, Laurie J. Hendren, and Alexandru Nicolau. Abstract description of pointer data structures: an approach for improving the analysis and optimization of imperative programs. ACM Letters on Programming Languages and Systems, 1(3):243–260, September 1992. CO-DEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http:

//www.acm.org/pubs/toc/Abstracts/
1057-4514/151644.html.

Pugh:1992:DDD

[19] William Pugh. Definitions of dependence distance. ACM Letters on Programming Languages and Systems, 1(3):261–265, September 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

URL http://www.acm.org/pubs/toc/
Abstracts/1057-4514/151645.html.

Ramkumar:1992:DLC

[20] Balkrishna Ramkumar. Distributed last call optimization for portable parallel logic programming. ACM Letters on Programming Languages and Systems, 1 (3):266–283, September 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/1057-4514/
151345.html.

Walker:1992:MIV

[21] Kenneth Walker and Ralph E. Griswold.

The maintenance of intermediate values in goal-directed evaluation. ACM Letters on Programming Languages and Systems, 1(3):284–298, September 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http://www.acm.

                                        org/pubs/toc/Abstracts/1057-4514/
                                        151341.html.

Fritzson:1992:GAD

[22] Peter Fritzson, Nahid Shahmehri, Mariam Kamkar, and Tibor Gyimothy. Generalized algorithmic debugging and testing. ACM Letters on Programming Languages and Systems, 1(4):303–322, December 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

                                        URL http://www.acm.org/pubs/toc/
                                        Abstracts/1057-4514/161498.html.

Landi:1992:USA

[23] William Landi. Undecidability of static analysis. ACM Letters on Programming Languages and Systems, 1(4):323–337, December 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

                                        URL http://www.acm.org/pubs/toc/
                                        Abstracts/1057-4514/161501.html.

Nilsen:1992:COS

[24] Kelvin D. Nilsen and William J. Schmidt. Cost-effective object space management for hardware-assisted real-time garbage collection. ACM Letters on Programming Languages and Systems, 1(4):338–354, December 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

                                        URL http://www.acm.org/pubs/toc/
                                        Abstracts/1057-4514/161508.html.

Srivastava:1992:UPO

[25] Amitabh Srivastava. Unreachable procedures in object-oriented programming. ACM Letters on Programming Languages and Systems, 1(4):355–364, December 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

                                        URL http://www.acm.org/pubs/toc/
                                        Abstracts/1057-4514/161517.html.

Ball:1993:WRC

[26] Thomas Ball. What’s in a region?: or computing control dependence regions in near-linear time for reducible control flow. ACM Letters on Programming Languages and Systems, 2(4):1–16, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

URL http://www.acm.org/pubs/toc/
Abstracts/1057-4514/176456.html.

Beaven:1993:ETE

[27] Mike Beaven and Ryan Stansifer. Explaining type errors in polymorphic languages. ACM Letters on Programming Languages and Systems, 2(4):17–30, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

URL http://www.acm.org/pubs/toc/
Abstracts/1057-4514/176460.html.

Binkley:1993:PEI

[28] David Binkley. Precise executable interprocedural slices. ACM Letters on Programming Languages and Systems, 2(4):31–45, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/1057-4514/
176473.html.

Boehm:1993:IML

[29] Hans-J. Boehm, Alan J. Demers, and Chris Uhler. Implementing multiple locks using Lamport’s mutual exclusion algorithm. ACM Letters on Programming Languages and Systems, 2(4):46–58, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

URL http://www.acm.org/pubs/toc/
Abstracts/1057-4514/176479.html.

Briggs:1993:ERS

[30] Preston Briggs and Linda Torczon. An efficient representation for sparse sets. ACM Letters on Programming Languages and Systems, 2(4):59–69, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

                                        URL http://www.acm.org/pubs/toc/
                                        Abstracts/1057-4514/176484.html.

Bumbulis:1993:RMV

[31] Peter Bumbulis and Donald D. Cowan. RE2C: a more versatile scanner generator. ACM Letters on Programming Languages and Systems, 2(4):70–84, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

                                        URL http://www.acm.org/pubs/toc/
                                        Abstracts/1057-4514/176487.html.

Cameron:1993:ECG

[32] Robert D. Cameron. Extending contextfree grammars with permutation phrases. ACM Letters on Programming Languages and Systems, 2(4):85–94, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

                                        URL http://www.acm.org/pubs/toc/
                                        Abstracts/1057-4514/176490.html.

Choudhary:1993:UCF

[33] Alok Choudhary, Geoffrey Fox, Seema Hiranandani, Ken Kennedy, Charles Koelbel, Sanjay Ranka, and Chau-Wen Tseng. Unified compilation of Fortran 77D and 90D. ACM Letters on Programming Languages and Systems, 2(4):95–114, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

                                        URL http://www.acm.org/pubs/toc/
                                        Abstracts/1057-4514/176495.html.

Day:1993:RRM

[34] Mark Day, Barbara Liskov, Umesh Maheshwari, and Andrew C. Myers. References to remote mobile objects in Thor. ACM Letters on Programming Languages and Systems, 2(4):115–126, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

URL http://www.acm.org/pubs/toc/
Abstracts/1057-4514/176500.html.

Eyre-Todd:1993:DDR

[35] Richard A. Eyre-Todd. The detection of dangling references in C++ programs. ACM Letters on Programming Languages and Systems, 2(4):127–134, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

URL http://www.acm.org/pubs/toc/
Abstracts/1057-4514/176504.html.

Gupta:1993:OAB

[36] Rajiv Gupta. Optimizing array bound checks using flow analysis. ACM Letters on Programming Languages and Systems, 2(4):135–150, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/1057-4514/
176507.html.

Kaser:1993:CID

[37] Owen Kaser, C. R. Ramakrishnan, and Shaunak Pawagi. On the conversion of indirect to direct recursion. ACM Letters on Programming Languages and Systems, 2(4):151–164, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/1057-4514/
176510.html.

Larus:1993:CSM

[38] James R. Larus. Compiling for sharedmemory and message-passing computers. ACM Letters on Programming Languages and Systems, 2(4):165–180, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

URL http://www.acm.org/pubs/toc/
Abstracts/1057-4514/176514.html.

Marriott:1993:PEG

[39] Kim Marriott and Harald Søndergaard. Precise and efficient groundness analysis for logic programs. ACM Letters on Programming Languages and Systems, 2(4):181–196, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/1057-4514/
176519.html.

Marriott:1993:SCL

[40] Kim Marriott and Peter J. Stuckey. Semantics of constraint logic programs with optimization. ACM Letters on Programming Languages and Systems, 2(4):197–212, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/1057-4514/
176522.html.

Metzger:1993:ICP

[41] Robert Metzger and Sean Stroud. Interprocedural constant propagation: an empirical study. ACM Letters on Programming Languages and Systems, 2 (4):213–232, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/1057-4514/
176526.html.

Qian:1993:RON

[42] Xiaolei Qian and Allen Goldberg. Referential opacity in nondeterministic data refinement. ACM Letters on Programming Languages and Systems, 2(4):233–241, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

URL http://www.acm.org/pubs/toc/
Abstracts/1057-4514/176578.html.